OpenStudy (anonymous):

Jason purchased a pack of game cards that was on sale for 13% off. The sales tax in his county is 8%. Let y represent the original price of the cards. Write an expression that can be used to determine the final cost of the cards.

OpenStudy (anonymous):

Jason purchased a pack of game cards that was on sale for 13% off. The sales tax in his county is 8%. Let y represent the original price of the cards. Write an expression that can be used to determine the final cost of the cards. y − 0.13y 1.08(0.87y) 0.08(0.87y) y − 0.87y + 0.08y

OpenStudy (anonymous):

y-0.13y y-0.87y+0.08y

OpenStudy (anonymous):

@SymbioteBowTie

OpenStudy (anonymous):

@J-bird

OpenStudy (anonymous):

PLEASE HELP IF YOU CAN!!!!

OpenStudy (anonymous):

I also have more :/

OpenStudy (anonymous):

I'd say last option

OpenStudy (anonymous):

ok thanks so much!!! i have 4 more can you help?

OpenStudy (anonymous):

I'll try

OpenStudy (anonymous):

Which expression is equivalent to one over fivem − 20? one over five(m − 4) one over five(m − 100) 5(m − 4) 5(m − 100)

OpenStudy (anonymous):

:/ its hard

OpenStudy (anonymous):

The first answer, it split the 20 into 5*4.

OpenStudy (anonymous):

ok thanks!!!!

OpenStudy (anonymous):

Pablo purchased adult and youth tickets for the football game. He bought x adult tickets for $74 each and y youth tickets for $35. Write an expression that can be used to show the total cost, C, of all tickets. C = (74 + 35)x C = (74 + 35)(x + y) C = 74y + 35x C = 74x + 35y Help on this one two?

OpenStudy (anonymous):

The last answer

OpenStudy (anonymous):

OK THANKSSS

OpenStudy (anonymous):

Write an expression for the perimeter of the triangle shown below: A triangle is shown with side lengths labeled 0.5x plus 9, 3x, and negative 3.4x minus 25. 6.9x − 16 6.9x + 16 0.1x − 16 0.1x + 16

OpenStudy (anonymous):

??

OpenStudy (anonymous):

6.9x-16

OpenStudy (anonymous):

0.1x-16

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

and the picture is the one with the perimiter

OpenStudy (anonymous):

oh that helps

OpenStudy (anonymous):

Write an expression that is equivalent to three over five (3y + 15). three over fivey + 9 three over five y + 15 nine over fivey + 9 nine over fivey + 15 This is the last one thanks so much for all the help if you answer this one ill give you a medal! I know you didnt finish the other one but so its ready xD

OpenStudy (anonymous):

wait i almost have the last one solved XD

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

for the last question, its the third answer 0.1x minus 16

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

and the other one is the last one xD

OpenStudy (anonymous):

ok, 9 over 5 plus 9, or the third answer

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!!!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

I gave you the best response!

OpenStudy (anonymous):

ill tell you what i get

OpenStudy (anonymous):

awesome ty!

OpenStudy (anonymous):

xD missed one sorry

OpenStudy (anonymous):

Jason purchased a pack of game cards that was on sale for 13% off. The sales tax in his county is 8%. Let y represent the original price of the cards. Write an expression that can be used to determine the final cost of the cards. y − 0.13y 1.08(0.87y) 0.08(0.87y) y − 0.87y + 0.08y

OpenStudy (anonymous):

0.0

OpenStudy (anonymous):

y- 0.13y y-0.87y+0.08y

OpenStudy (anonymous):

lol sorry if you can answer that then i can submit it lol

OpenStudy (anonymous):

The last answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok it was a 80%

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

4 of 5 sorry

OpenStudy (anonymous):

thats fine

OpenStudy (anonymous):

xD thanks so much though

OpenStudy (anonymous):

np

OpenStudy (anonymous):

alrighty see ya:D

OpenStudy (anonymous):

you too

OpenStudy (anonymous):

and thx to you, i got my smart score to 50!